A continuous random variable x has the pdf given by

f(x)={k(1-x) for 0<x<1 both inclusive

{o,elsewhere

find t,the value of the constant k. Also find the mean and the variance of x


1
Expert's answer
2021-07-15T11:25:51-0400
"\displaystyle\int_{-\infin}^{\infin}f(x)x=\displaystyle\int_{0}^{1}k(1-x)dx"

"=k[x-\dfrac{x^2}{2}]\begin{matrix}
 1 \\
 0
\end{matrix}=\dfrac{1}{2}k=1"

"k=2"

"mean=\mu=E[X]=\displaystyle\int_{-\infin}^{\infin}xf(x)dx"

"=\displaystyle\int_{0}^{1}2x(1-x)dx=[x-\dfrac{2x^3}{3}]\begin{matrix}
 1\\
 0
\end{matrix}=\dfrac{1}{3}"


"E[X^2]=\displaystyle\int_{-\infin}^{\infin}x^2f(x)dx=\displaystyle\int_{0}^{1}2x^2(1-x)dx"

"=[\dfrac{2x^3}{3}-\dfrac{x^4}{2}]\begin{matrix}
 1 \\
 0
\end{matrix}=\dfrac{1}{6}"


"Var(X)=E[X^2]-(E[X])^2"

"=\dfrac{1}{6}-(\dfrac{1}{3})^2=\dfrac{1}{18}"